OpenStudy (anonymous):

What is the equation of a circle whose center is at the origin and whose radius is 16?

OpenStudy (apoorvk):

For a circle with center at (a,b) and a radius 'r', it's equation is: \[(x-a)^2 + (y-b)^2 = r^2\] What do you think now?

OpenStudy (anonymous):

(x-a)^2+(y-b)^2=256??